OpenStudy (sharkfinsoup):

what is the local maximum value of the function g(x)=x^3+5x^2-17x-21

OpenStudy (sharkfinsoup):

I know I have to find the x value first and then plug it into g(x) but I'm not sure how to get the x value.

OpenStudy (irishboy123):

you doing calculus or plotting graphs ?!?!

OpenStudy (sharkfinsoup):

precalc

OpenStudy (irishboy123):

assume that means you are not allowed to differentiate but your thing factorises.... \((x + 1) (x - 3) (x + 7)\) and looks like this....  so that might even help

OpenStudy (sharkfinsoup):

But then how do I take that information and use it to find the local maximum value

OpenStudy (irishboy123):

you know what....right now i haven't a clue how to help you without using calculus :(

OpenStudy (sharkfinsoup):

Use the calculus I'm willing to try it

OpenStudy (inkyvoyd):

here is an interesting way

OpenStudy (inkyvoyd):

shift your graph up until you get a double root, then you have found your minimum

OpenStudy (inkyvoyd):

you want your root at -1 and 3 to coalesce

OpenStudy (irishboy123):

Hang on @sharkfinsoup, smart dude at work !!

OpenStudy (sharkfinsoup):

ok take your time @IrishBoy123

OpenStudy (sharkfinsoup):

What do I do after I shift the graph? @inkyvoyd

OpenStudy (inkyvoyd):

g(x)=x^3+5x^2-17x-21 g(x)+a=x^3+5x^2-17x-21+a=(x+7)(x-b)^2 so \(x^3+5x^2-17x-21+a=(x+7)(x-b)^2\) \(x^3+5x^2-17x-21+a=7 b^2 - 14 b x + b^2 x + 7 x^2 - 2 b x^2 + x^3\) so let's try to solve for a and b...

OpenStudy (inkyvoyd):

oops made a mistake... we don't know the root will stay at x=-7 unfortunately

OpenStudy (inkyvoyd):

so actually, \(g(x)+a=x^3+5x^2-17x-21+a=(x+b)(x-c)^2\)

OpenStudy (inkyvoyd):

so \(x^3+5x^2-17x-21+a=b c^2 - 2 b c x + b x^2 + c^2 x - 2 c x^2 + x^3\) Now simplify: \(5x^2-17x-21+a=b c^2 - 2 b c x + b x^2 + c^2 x - 2 c x^2\)

OpenStudy (inkyvoyd):

Group terms: \(0=(b-2c-5)x^2+(17-2bc+c^2)x+(21-a+bc^2)\)

OpenStudy (inkyvoyd):

oh gosh darnit all my progress goes toward finding the local minimum LOL... well that is a good thing in the end since you can use my method to figure out the local max on your own

OpenStudy (inkyvoyd):

So LHS=0=RHS=0 so all coefficients must be 0... We need to solve a yucky system of 3 equations in 3 unknowns with quadratics, but it's doable

OpenStudy (inkyvoyd):

0=b-2c-5 0=17-2bc+c^2 0=21-a+bc^2

OpenStudy (inkyvoyd):

if you do a ton of algebra you will get http://prntscr.com/di06yr

OpenStudy (inkyvoyd):

so our minima must be at x=a = \(-\frac{16}{27}(28+19\sqrt{19})\) or something of the sort...

OpenStudy (unklerhaukus):

that's not a nice result.

OpenStudy (unklerhaukus):

and it doesn't agree with the plot.

OpenStudy (inkyvoyd):

might be 28-19sqrt(19)

OpenStudy (inkyvoyd):

\(-\frac{16}{27}(28-19\sqrt{19})\) there we go, your local minima

OpenStudy (inkyvoyd):

FYI @UnkleRhaukus I did the calculus and it is not a nice result. it IS my result though so I'm happy.

OpenStudy (inkyvoyd):

but fyi sharkfin your numbers are so nasty that I wouldn't want to do this problem if calculus WAS allowed.

OpenStudy (unklerhaukus):

you had \(x = −\frac{16}{27}(28+19\sqrt{19}) \), i think you want \(y = \frac{16}{27}(28+19\sqrt{19}) \)

OpenStudy (sharkfinsoup):

Alright well for anyone curious the correct answer was 65.67